ELUCIDATION OF EXTRA-MECHANICAL EFFECTS IN REDUCTION OF INFRARED AND RAMAN SPECTRA OF DIATOMIC MOLECULES IN GASEOUS SAMPLES

K. L. BAK, UNI-C, Olof Palmes Alle 38, DK-8200 Aarhus N, Denmark; S. P. A. SAUER, Chemistry Laboratory IV, Department of Chemistry, University of Copenhagen, DK-2100 Copenhagen Ø, Denmark; J. ODDERSHEDE, University of Southern Denmark, Department of Chemistry, Campusvej55, DK-5230 Odense M, Denmark; \underline{J. F. OGILVIE}, Escuela de Quimica, Universidad de Costa Rica, San Pedro, San Jose, Costa Rica, and Centre for Experimental and Constructive Mathematics, Department of Mathematics, Simon Fraser University, 8888 University Drive, Burnaby, British Columbia V5A 1S6 Canada.